Circular No. 7033

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998eg IN UGC 12133 G. M. Hurst, Basingstoke, England, reports the discovery by Tom

Boles, Wellingborough, of a supernova (mag 16.0) on 80-s unfiltered CCD images obtained on Oct. 19.906 UT for the U.K. Supernova Patrol with a 0.26-m reflector; it was not visible on Sept. 17 (limiting mag 18). Boles provides the precise position α = 22^{h}39^{m}30^{s}.26, δ = +8^{o}36′21′′.3 (equinox 2000.0). Additional reported CCD magnitudes (unfiltered unless otherwise noted): Oct. 19.896, 15.7 (M. Armstrong, Rolvenden, England); 20.88, R ∼ 16.4 (S. Benetti, F. Ghinassi, E. Marchetti, G. Tessicini, C. Vuerli, and A. Zacchei, La Palma); 20.995, 15.5 (Armstrong); 21.25, R = 17.6 (D. Balam, University of Victoria); 21.385, 16.2 (Y. Sano, Nayoro, Hokkaido, Japan). Additional reported position end figures: 30^{s}.37, 21′′.0 (Armstrong, 0.26-m reflector); 30^{s}.29, 20′′.6 (Balam, 1.82-m Plaskett telescope); 30^{s}.35, 20′′.5 (Sano, 0.28-m reflector; communicated by H. Yamaoka, Kyushu University). Balam gives the supernova’s offset as 25′′.9 west and 25′′.3 south of the nucleus of UGC 12133, and notes that there is nearby star (R = 15.8) at position end figures 30^{s}.08, 37′22′′.7; Balam has posted an image of SN

1998eg at http://astrowww.phys.uvic.ca/~balam/sn97do.html.

COMET C/1998 U1 (LINEAR) MPEC 1998-U12 provides detailed information concerning yet another

comet discovered by LINEAR. The object, discovered on Oct. 18, was reported as asteroidal and fast-moving, and a request for further observations was placed on The NEO Confirmation Page. From the confirmatory observations, made at four observatories on Oct. 20 and 21, it quickly became clear that the object had a retrograde, nearly parabolic orbit. Referring to the observations made with the 0.65-m f/3.6 reflector at the Ondřejov Observatory on Oct. 20.9 UT, P. Pravec notes that a faint tail was visible on individual 3-min exposures with a clear filter. The tail, extending 16′′ in p.a. 60^{o}, was confirmed on co-addition of six frames. The discovery observation and orbital elements follow:

1998 UT  α_{2000}  δ_{2000}  m_{1}
Oct. 18.40062  3^{h}56^{m}57.^{s}57  +33^{o}36′22′′.5  19.2
T = 1998 June 1.479 TT  ω = 130.^{o}116
Ω = 210.409 }2000.0
q = 4.06998 AU  i = 156.468
1998 October 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green